Presented below is selected information related to Bonita Electronics inventory. HDTV 42 HDTV- TV 50 Stand (per unit) $370
Answer #1
NRV LCNRV
HDTV 50 (474-27-46) = 401 370
HDTV 42 (290-13-21) = 256 256
TV stand (70-2-6) = 62 53

Comment if you face any issues